# TAIL: Text-Audio Incremental Learning

### v1

@@ Abstract
1 Introduction
@@
Multimodal interaction has always been a key mode of communication in the human world, particularly with the rise of language assistants (e.g. ChatGPT [40]). Understanding the modalities between text and audio has become a prominent area of research. Among these, the text-audio retrieval task is widely applied, aiming to search an audio clip or caption within a dataset based on a query (text or audio clip) from a different modality.
Figure 1: Comparison of finetuning (left) and our proposed prompt tuning method (right) for incremental learning. In incremental learning, finetuning updates all parameters, while we train only a subset.
@@
Inspired by advancements in Large Language Models (LLM), pre-trained models on multimodal data have achieved remarkable capabilities in multimodal understanding. However, with the rapid growth of data scale and the diversity of datasets, text-audio retrieval tasks still encounter two challenges in real-world applications, shown in Fig. 1. (1) Resource-Intensive Training. Training with large-scale datasets increases the model’s parameter requirements, thereby consuming substantial computational resources. (2) Catastrophic Forgetting. The introduction of new datasets (e.g. Dataset M) often disrupts the model’s previously learned distributions from previous old datasets (e.g. Dataset 1,…,M-1), resulting in catastrophic forgetting.
@@
There are many research works focusing on these challenges. (1) To address the problem of parameter efficiency. In recent years, prompt tuning methods, initially emerging in the field of parameter-efficient transfer learning, have also proven effective in incremental learning. In the image domain, methods such as those in [59, 53, 27, 57, 45, 58, 48, 34, 12, 65, 11] train only a small set of parameters, specifically prompts, to guide a pretrained model in incremental learning tasks without requiring storage of past examples. In the video domain, approaches such as [69, 42, 52] utilize local and global prompts to focus on temporal and spatial features. Prompt tuning demonstrates strong potential for mitigating catastrophic forgetting, as it keeps the backbone fixed and learns only prompts, significantly reducing training costs and improving learning efficiency. Despite its success in visual domains, this new paradigm remains largely unexplored in the text-audio learning domain. (2) To address the problem of catastrophic forgetting, various incremental learning approaches have been developed, including knowledge distillation based [1, 73, 16], exemplar-based [44, 24], representation-based [67, 18, 37, 32] and regularization-based [21, 31, 4, 46, 19] methods. These approaches primarily target tasks in continual or incremental video classification [43, 73, 41], image classification [14, 49], language or joint-vision-language tasks [26, 47], and self-supervised representation learning and pretraining [13, 4, 32]. However, these methods often encounter limitations: some achieve suboptimal performance due to restricted buffer sizes, while others require retraining all model parameters, resulting in increased training costs and substantial resource demands. Furthermore, these methods rarely conduct research specifically focused on text-audio learning.
Figure 2: Comparison of our method with previous methods in terms of forgetting degree (x-axis), performance (y-axis), and parameter scale (circular shadow). Finetuning suffers from severe forgetting issues, which our approach successfully addresses in a parameter-efficient way.
@@
In this paper, we introduce a novel task called Text-Audio Incremental Learning (TAIL). TAIL is aimed at training models to perform text-audio retrieval continuously across new datasets without forgetting knowledge from previous ones. For this task, we propose a parameter-efficient method Prompt Tuning for Audio-Text Incremental Learning (coined as PTAT) as a starting point to benchmark the TAIL task. In PTAT, we propose two key modules. For the resource-intensive training challenge, we proposed a parameter-efficient Audio-Text Prompt Generation (ATPG) module to efficiently bridge the gap between the audio and text features and learn new knowledge. Moreover, for the catastrophic forgetting challenge, to preserve and leverage the previously learned knowledge, we design an Audio-Text Similarity and Feature Distillation (ATSFD) module to distill the learned audio feature, text feature, and their similarity into new incremental steps, which enables the model to preserve previously learned cross-modal audio-text correlations in new datasets. As shown in Fig. 2, we calculate the forgetting level based on the percentage drop in average $R@10$. Compared to other methods, our proposed PTAT model achieves superior retrieval performance with fewer parameters and retains the lowest model forgetting level in text-audio incremental learning across all datasets.
@@
  • We developed the Prompt Tuning for Audio-Text incremental learning (PTAT) method that achieves efficient performance with a minimal number of parameters, effectively alleviating the issue of catastrophic forgetting on previous datasets as new datasets are incrementally added.
@@
  • We experimentally demonstrate that PTAT outperforms the state-of-the-art on four audio-text datasets including AudioCaps, Clotho, BBC Sound Effects, and Audioset.
2 Related work
2.1 Text-to-Audio Retrieval
@@
The Text-to-Audio retrieval task [62, 63, 39], which aims to retrieve relevant text from a text database based on a given audio input, has garnered considerable attention. DiffATR [62] introduces a diffusion-based generative framework for Audio-Text Retrieval, modeling the joint probability distribution between text and audio, and demonstrating effective retrieval performance and adaptability to out-of-domain samples on the AudioCaps and Clotho datasets. In contrast, recent work [39] addresses the under-explored issue of temporal order comprehension in audio, evaluating this capability using the AudioCaps and Clotho datasets, introducing a synthetic dataset named SynCaps, and proposing a text-text contrastive loss to enhance retrieval accuracy. Additionally, LGMM [55] presents a local-to-global multiscale matching method that refines multiscale alignment by leveraging intrinsic relative correlations within each modality, thus effectively addressing challenges associated with binary cross-modal matching labels. However, these works have ignored the generalizability of audio/text retrieval tasks to new datasets. Our proposed method can maintain the retrieval ability of the model itself while adapting to new datasets.
2.2 Audio Incremental Learning
@@
Incremental or continual learning [54, 36, 70, 71] is defined as training a learner for several tasks sequentially, without forgetting knowledge obtained from the preceding tasks. The data from the old tasks is no longer available when training new tasks. Various methods have been proposed to prevent catastrophic forgetting in continual learning. DFWF [30] uses a knowledge distillation loss to preserve memory from the original model, and static memory networks [25] introduce static memory to reduce memory usage and model complexity. Few-shot CL [56] enables fast and interactive model updates in a few-shot learning framework to expand the audio classifier to recognize novel classes, while AI-CL [38] uses a combination of the parameter-efficient Convolutional Adapter and the compute-efficient Frequency-Time factorized Attention to efficiently learns new tasks and prevents catastrophic forgetting. IODFD [36] propose a cosine similarity-based distillation loss and use it along with a Kullback-Leibler divergence-based distillation loss to preserve knowledge about the old classes. Currently, there has been no work focused on Audio-Text incremental learning. In this paper, we introduce a text-audio incremental learning task to address this missing task in practical applications.
2.3 Prompt-based Transfer Learning
@@
Many recent works have focused on efficient transfer learning and fine-tuning techniques for downstream tasks, especially using the prompt tuning method, a highly lightweight and effective method for enabling model transfer across different tasks or datasets. These methods achieve efficient fine-tuning by inserting learnable prompt vectors at different locations inside a transformer encoder while freezing other parameters during training.  [20, 17] proposed an effective and efficient visual prompt tuning approach for large-scale transformer-based model adaptation. PETL_AST [3] presents a detailed investigation of common parameter-efficient methods including prompt tuning, adapters, and LoRA on Audio Spectrogram Transformer. For cross-modal learning, I-VL [23] uses text prompts to bridge the gap between static images and videos. [5, 29, 60, 50] added learnable prompts to the input of the text encoder to enhance the model’s learning capacity on new tasks.  [69, 66] used interactive prompts across different modalities to fine-tune both uni-modal and multi-modal features.  [68, 2, 7, 64] achieved a significant accuracy improvement when adapting CLIP to unseen domains with only training a lightweight prompt generator. Our approach PTAT is based on the state-of-the-art model [66] to perform parameter optimization.
3 Method
@@
Figure 3: Overview of our framework: Our model framework comprises two key modules: Audio-Text Prompt Generation (ATPG) and Audio-Text Similarity and Feature Distillation (ATSFD). (a) Within ATPG, we generate a set of interconnected, trainable global audio prompts, denoted as $\mathcal{A}$ (purple), along with prefix (light blue) and postfix (dark blue) text prompts, $\mathcal{T}_{pre}$ and $\mathcal{T}_{post}$. During training, both the audio and text encoders are kept frozen, with prompts added to facilitate optimization. (b) In ATSFD, we apply feature distillation and similarity distillation between the current step $m$ and the previous step $m-1$ to retain model knowledge. The gray line represents the similarity distillation loss $\mathcal{L}_{SD}$, while the green and orange lines denote text feature distillation and audio feature distillation in $\mathcal{L}_{FD}$.
@@ 3.2 Overall Framework
Audio-Text Prompt Generation (ATPG). Inspired by the prompt-based incremental learning paradigm in the vision domain [66] to propose an effective audio-text joint prompting approach. Mel spectrogram can retain the original audio sequence characteristics by splitting it into patches and adding position encoding.
@@
For text, inspired by [22], we implement a set of deep text prompts to enhance text representation, consisting of prefix text prompts (light blue)) $\mathcal{T}_{pre}=\{\mathcal{T}^{j}_{pre}\in\mathbb{R}^{d}\mid j\in\mathbb{N},1\leq j\leq n_{pre}^{t}\}$ and postfix text prompts (dark blue) $\mathcal{T}_{post}=\{\mathcal{T}^{j}_{post}\in\mathbb{R}^{d}\mid j\in\mathbb{N},1\leq j\leq n_{post}^{t}\}$. The prefix text prompts are prepended to the text query before word embedding, while the postfix text prompts are appended afterward. Here, $d$ is the dimensionality of the prompt vectors, and $n_{pre}^{t}$ and $n_{post}^{t}$ represent the numbers of prefix and postfix prompts, respectively. The generated text prompts are input to the first layer of the text encoder, where they are append/prepend to the text embeddings.
The audio and text prompts mentioned above are interconnected. As shown in Fig. 3 ATPG module, we design a simple and efficient prompt that can effectively connect two modalities. Inspired by  [66], we consider using two simple linear layers, $S^{pre}_{linear}$ and $S^{post}_{linear}$, to map the audio prompts $\mathcal{A}$ to the text prefix and text postfix prompts. This process can be formulated as follows:
$\mathcal{T}_{pre}$ $=S^{pre}_{linear}(\mathcal{A}),$ (1)
$\mathcal{T}_{post}$ $=S^{post}_{linear}(\mathcal{A})$
@@
Audio-Text Similarity and Feature Distillation (ATSFD). A crucial aspect of incremental learning is preserving cross-modal semantic similarity to mitigate forgetting when learning new data [51, 35]. To achieve this, as shown in Fig. 3, we designed the Audio-Text Similarity and Feature Distillation module.
@@ 3.3 Training Objectives
$\ell_{t2a}$ $=\frac{1}{N}\sum_{i=0}^{N}\log\text{diag}(\text{softmax}(C_{t2a})),$ (6)
$\ell_{a2t}$ $=\frac{1}{N}\sum_{i=0}^{N}\log\text{diag}(\text{softmax}(C_{a2t}))$
@@
$\mathcal{L}_{\text{contrast}}=\ell_{t2a}+\ell_{a2t}$ (7)
Furthermore, we apply $KL$ divergence to align the feature distributions of the output from the audio encoder and text encoder, enhancing the accuracy of audio-text retrieval. This is expressed as:
$\mathcal{L}_{\text{kl}}=\mathbb{E}_{(X^{a},X^{t})\sim\mathcal{D}_{k}}\left[\text{KL}\left(E_{a}\parallel E_{t}\right)+\text{KL}\left(E_{t}\parallel E_{a}\right)\right]$ (8)
We propose two additional losses to address incremental learning tasks. To mitigate the forgetting issue in incremental learning and ensure consistency and alignment of the model across different feature spaces, we apply the losses $\mathcal{L}_{FD}$ and $\mathcal{L}_{SD}$ defined in Equ. 3 and Equ. 5.
@@
$\mathcal{L}=\mathcal{L}_{\text{kl}}+\lambda\mathcal{L}_{\text{contrast}}+\mathcal{L}_{\text{FD}}+\alpha\mathcal{L}_{\text{SD}}$ (9)
@@
4 Experiments
4.1 Experimental Setup
Datasets. We conduct experiments with our PTAT framework, comparing it to baselines on the AudioCaps [28], Clotho [8], AudioSet SL, and BBC Sound Effects subsets within the WavCaps [33] dataset. AudioCaps [28] contains approximately 50,000 audio clips sourced from AudioSet [15], the largest audio event dataset, and is annotated by human listeners. Clotho [8] comprises around 6,000 audio clips from the FreeSound2 platform, each clip accompanied by five human-annotated captions. The AudioSet SL and BBC Sound Effects subsets are part of the WavCaps [33] dataset.
Since neither AudioSet SL nor BBC Sound Effects provides predefined training and test splits, we manually divided the data for these datasets. The AudioSet SL subset contains 58,000 audio clips (after excluding portions overlapping with AudioCaps from the original AudioSet SL), which we randomly split into a training set of 55,000 samples and a test set of 3,000 samples. The BBC Sound Effects dataset consists of more than 30,000 audio clips recorded worldwide over the past century, which we randomly divided into a training set of 28,000 samples and a test set of 2,000 samples.
For our incremental learning setup, we define four incremental steps: AudioCaps as the first step, Clotho as the second, BBC Sound Effects as the third, and AudioSet SL as the fourth.
Table 1: Comparison with Baselines: Results on Each Dataset After Sequential Training Across Four Datasets(AudioCaps, Clotho, BBC Sound Effects, and Audioset). The results present the model’s performance at the Current Step of the experiment.
Method Parameters AudioCaps (Trained on Step 1) Clotho (Trained on Step 2)

### v2

@@ Abstract.
Text-Audio Retrieval, Prompt Tuning, Incremental Learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Natural language processingccs: Computing methodologies Speech recognition
1. Introduction
Figure 1. An example of Step-wise Streaming of Text-Audio Datasets. Each dataset’s modality distribution is different. At each step, the model is trained on an unseen dataset, and tested on historical datasets.
@@ 1. Introduction
Multimodal interaction has always been a key mode of communication in the human world, particularly with the rise of language assistants (e.g. ChatGPT (OpenAI, 2022)). Understanding the modalities between text and audio has become a prominent area of research. Among these, the text-audio retrieval task is widely applied, aiming to search an audio clip or caption within a dataset based on a query (text or audio clip) from a different modality.
@@
Inspired by advancements in Large Language Models (LLM), pre-trained models on multimodal data have achieved remarkable capabilities in multimodal understanding. However, with the rapid growth of data scale and the diversity of datasets, text-audio retrieval tasks still encounter two challenges in real-world applications, (1) Resource-Intensive Training. Training with large-scale datasets increases the model’s parameter requirements, thereby consuming substantial computational resources. (2) Catastrophic Forgetting. The introduction of unseen datasets often disrupts the model’s previously learned distributions from previous old datasets, resulting in catastrophic forgetting.
@@
There are many research works focusing on these challenges. (1) To address the problem of parameter efficiency. In recent years, prompt tuning methods, initially emerging in the field of parameter-efficient transfer learning, have also proven effective in incremental learning. In the image domain, methods such as those in (Wang et al., 2022c, 2024b; Khan et al., 2023; Wang et al., 2022a; Smith et al., 2023; Wang et al., 2022b; Tang et al., 2023a; Menabue et al., 2024; Feng et al., 2024b; Yang et al., 2024a; Feng et al., 2024a) train only a small set of parameters, specifically prompts, to guide a pretrained model in incremental learning tasks without requiring storage of past examples. In the video domain, approaches such as (Zhao et al., 2024; Pei et al., 2023a; Villa et al., 2023) utilize local and global prompts to focus on temporal and spatial features. Prompt tuning demonstrates strong potential for mitigating catastrophic forgetting, as it keeps the backbone fixed and learns only prompts, significantly reducing training costs and improving learning efficiency. Despite its success in visual domains, this new paradigm remains largely unexplored in the text-audio learning domain.(2) To address the problem of catastrophic forgetting, various incremental learning approaches have been developed, including knowledge distillation-based (Boschini et al., 2022; Zuo et al., 2024; Guo et al., 2024), exemplar-based (Rebuffi et al., 2017; Kang et al., 2022), representation-based (Zhang et al., 2023b; Huang et al., 2025; Mushtaq et al., 2024; Marczak et al., 2024)and regularization-based (Joseph et al., 2022; Ma et al., 2024; Cha et al., 2024; Song et al., 2023; Huang et al., [n. d.]) methods. These approaches primarily target tasks in continual or incremental video classification (Pian et al., 2023; Zuo et al., 2024; Pei et al., 2023b), image classification (Fu et al., 2024; Tang et al., 2023b), language or joint-vision-language tasks (Ke et al., 2023; Srinivasan et al., 2022), and self-supervised representation learning and pretraining (Fini et al., 2021; Cha et al., 2024; Marczak et al., 2024). However, these methods often encounter limitations: some achieve suboptimal performance due to restricted buffer sizes, while others require retraining all model parameters, resulting in increased training costs and substantial resource demands. Furthermore, these methods rarely conduct research specifically focused on text-audio learning.
Figure 2. Comparison of our method with previous methods in terms of forgetting degree (x-axis), performance (y-axis), and parameter scale (circular shadow). Experimental results demonstrate that our method achieves the best retrieval performance and the strongest resistance to forgetting, while maintaining the highest parameter efficiency.
@@
In this paper, we introduce a novel multimodal task called Text-Audio Incremental Learning (TAIL), where the model is progressively updated with new datasets at each step, as shown in Fig. 1. TAIL is aimed at training models to perform text-audio retrieval continuously across new datasets without forgetting knowledge from previous ones. For this task, we propose a parameter-efficient method Prompt Tuning for Audio-Text Incremental Learning (PTAT) as a starting point to benchmark the TAIL task. In PTAT, we propose two key modules. For the resource-intensive training challenge, we proposed a parameter-efficient Audio-Text Prompt Generation (ATPG) module to efficiently bridge the gap between the audio and text features and learn new knowledge. Moreover, for the catastrophic forgetting challenge, to preserve and leverage the previously learned knowledge, we design an Audio-Text Similarity and Feature Distillation (ATSFD) module to distill the learned audio feature, text feature, and their similarity into new incremental steps, which enables the model to preserve previously learned cross-modal audio-text correlations in new datasets. To intuitively assess the model’s ability to cope with catastrophic forgetting, we introduce an evaluation metric termed the Anti-Forgetting Score. This metric measures the model’s capacity to retain early-learned knowledge by comparing its performance on the same dataset before and after incremental learning. As illustrated in the Fig. 2, we select the AudioCaps dataset, used in the first stage of training, as the evaluation benchmark to clearly demonstrate the model’s forgetting resistance. Compared to other state-of-the-arts methods, our proposed PTAT model achieves superior retrieval performance with fewer parameters and retains the lowest model forgetting level in text-audio incremental learning across all datasets.
@@
  • We propose a parameter-efficient method, Prompt Tuning for Audio-Text Incremental Learning (PTAT), which introduces an Audio-Text Prompt Generation (ATPG) module for efficient cross-modal representation learning, and an Audio-Text Similarity and Feature Distillation (ATSFD) module to mitigate catastrophic forgetting by preserving audio-text correlations across incremental steps.
@@
  • Extensive experiments demonstrate that PTAT outperforms state-of-the-art methods on four benchmark audio-text datasets (AudioCaps, Clotho, BBC Sound Effects, and AudioSet), achieving a 4.46% performance gain while using only 2.42% of the parameters compared to the full-parameter Finetune (Sequential) method.
2. Related work
2.1. Text-to-Audio Retrieval
@@
The Text-to-Audio retrieval task (Xin et al., 2024a, b; Oncescu et al., [n. d.]), which aims to retrieve relevant text from a text database based on a given audio input, has garnered considerable attention. DiffATR (Xin et al., 2024a) introduces a diffusion-based generative framework for Audio-Text Retrieval, modeling the joint probability distribution between text and audio, and demonstrating effective retrieval performance and adaptability to out-of-domain samples on the AudioCaps and Clotho datasets. In contrast, recent work (Oncescu et al., [n. d.]) addresses the under-explored issue of temporal order comprehension in audio, evaluating this capability using the AudioCaps and Clotho datasets, introducing a synthetic dataset named SynCaps, and proposing a text-text contrastive loss to enhance retrieval accuracy. Additionally, LGMM (Wang et al., 2024a) presents a local-to-global multiscale matching method that refines multiscale alignment by leveraging intrinsic relative correlations within each modality, thus effectively addressing challenges associated with binary cross-modal matching labels. However, these works have ignored the generalizability of audio/text retrieval tasks to new datasets. Our proposed method can maintain the retrieval ability of the model itself while adapting to new datasets.
2.2. Audio Incremental Learning
@@
Incremental or continual learning (Wang et al., 2024c; Mulimani and Mesaros, 2024; Zhou et al., 2024a, b) is defined as training a learner for several tasks sequentially, without forgetting knowledge obtained from the preceding tasks. The data from the old tasks is no longer available when training new tasks. Various methods have been proposed to prevent catastrophic forgetting in continual learning. DFWF (Ma et al., 2021) uses a knowledge distillation loss to preserve memory from the original model, and static memory networks (Karam et al., 2022) introduce static memory to reduce memory usage and model complexity. Few-shot CL (Wang et al., 2021) enables fast and interactive model updates in a few-shot learning framework to expand the audio classifier to recognize novel classes, while AI-CL (Muthuchamy Selvaraj et al., 2023) uses a combination of the parameter-efficient Convolutional Adapter and the compute-efficient Frequency-Time factorized Attention to efficiently learns new tasks and prevents catastrophic forgetting. IODFD (Mulimani and Mesaros, 2024) propose a cosine similarity-based distillation loss and use it along with a Kullback-Leibler divergence-based distillation loss to preserve knowledge about the old classes. Currently, there has been no work focused on Audio-Text incremental learning. In this paper, we introduce a text-audio incremental learning task to address this missing task in practical applications.
2.3. Prompt-based Transfer Learning
@@
Many recent works have focused on efficient transfer learning and fine-tuning techniques for downstream tasks, especially using the prompt tuning method, a highly lightweight and effective method for enabling model transfer across different tasks or datasets. These methods achieve efficient fine-tuning by inserting learnable prompt vectors at different locations inside a transformer encoder while freezing other parameters during training.  (Jia et al., 2022; Han et al., 2023) proposed an effective and efficient visual prompt tuning approach for large-scale transformer-based model adaptation. PETL_AST (Cappellazzo et al., 2024) presents a detailed investigation of common parameter-efficient methods including prompt tuning, adapters, and LoRA on Audio Spectrogram Transformer. For cross-modal learning, I-VL (Ju et al., 2022a) uses text prompts to bridge the gap between static images and videos. (Chen et al., [n. d.]; Li et al., 2024; Wu et al., 2023b; Tian et al., 2024) added learnable prompts to the input of the text encoder to enhance the model’s learning capacity on new tasks. (Zhao et al., 2024; Yang et al., 2024b) used interactive prompts across different modalities to fine-tune both uni-modal and multi-modal features. (Zhang et al., 2023a; Cao et al., 2025; Deng et al., 2023; Xue et al., 2023) achieved a significant accuracy improvement when adapting CLIP to unseen domains with only training a lightweight prompt generator. Our approach PTAT is based on the state-of-the-art model (Yang et al., 2024b) to perform parameter optimization.
3. Method
@@
Figure 3. Overview of our framework. Our model consists of two modules: Audio-Text Prompt Generation (ATPG) and Audio-Text Similarity and Feature Distillation (ATSFD). (a) Within ATPG, we generate a set of interconnected, trainable global audio prompts, denoted as $\mathcal{A}$ (purple), along with prefix (light blue) and postfix (dark blue) text prompts, $\mathcal{T}_{pre}$ and $\mathcal{T}_{post}$. During training, both the audio and text encoders are kept frozen, with prompts added to facilitate optimization. (b) In ATSFD, we apply feature distillation and similarity distillation between the current Step $m$ and the previous Step $m-1$ to retain model knowledge. The gray line represents the similarity distillation loss $\mathcal{L}_{SD}$, while the green and orange lines denote text feature distillation and audio feature distillation in $\mathcal{L}_{FD}$.
@@ 3.2. Overall Framework
Audio-Text Prompt Generation (ATPG). Inspired by the prompt-based incremental learning paradigm in the vision domain (Yang et al., 2024b) to propose an effective audio-text joint prompting approach. Mel spectrogram can retain the original audio sequence characteristics by splitting it into patches and adding position encoding.
@@
For text, inspired by (Ju et al., 2022b), we implement a set of deep text prompts to enhance text representation, consisting of prefix text prompts (light blue)) $\mathcal{T}_{pre}=\{\mathcal{T}^{j}_{pre}\in\mathbb{R}^{d}\mid j\in\mathbb{N},1\leq j\leq n_{pre}^{t}\}$ and postfix text prompts (dark blue) $\mathcal{T}_{post}=\{\mathcal{T}^{j}_{post}\in\mathbb{R}^{d}\mid j\in\mathbb{N},1\leq j\leq n_{post}^{t}\}$. The prefix text prompts are prepended to the text query before word embedding, while the postfix text prompts are appended afterward. Here, $d$ is the dimensionality of the prompt vectors, and $n_{pre}^{t}$ and $n_{post}^{t}$ represent the numbers of prefix and postfix prompts, respectively. The generated text prompts are input to the first layer of the text encoder, where they are append/prepend to the text embeddings.
The audio and text prompts mentioned above are interconnected. As shown in Fig. 3 ATPG module, we design a simple and efficient prompt that can effectively connect two modalities. Inspired by  (Yang et al., 2024b), we consider using two simple linear layers, $S^{pre}_{linear}$ and $S^{post}_{linear}$, to map the audio prompts $\mathcal{A}$ to the text prefix and text postfix prompts. This process can be formulated as follows:
(1) $\mathcal{T}_{pre}$ $=S^{pre}_{linear}(\mathcal{A}),$
$\mathcal{T}_{post}$ $=S^{post}_{linear}(\mathcal{A}),$
@@
Audio-Text Similarity and Feature Distillation (ATSFD). A crucial aspect of incremental learning is preserving cross-modal semantic similarity to mitigate forgetting when learning new data (Tian and Xu, 2021; Mo and Tian, 2022). To achieve this, as shown in Fig. 3, we designed the Audio-Text Similarity and Feature Distillation module.
@@ 3.3. Training Objectives
(6) $\mathcal{L}_{t2a}$ $=\frac{1}{N}\sum_{i=0}^{N}\log\text{diag}(\text{softmax}(C_{t2a})),$
$\mathcal{L}_{a2t}$ $=\frac{1}{N}\sum_{i=0}^{N}\log\text{diag}(\text{softmax}(C_{a2t})).$
@@
(7) $\mathcal{L}_{\text{contrast}}=\mathcal{L}_{t2a}+\mathcal{L}_{a2t}.$
Furthermore, we apply $KL$ divergence to align the feature distributions of the output from the audio encoder and text encoder, enhancing the accuracy of audio-to-texttext-to-audio retrieval. This is expressed as:
(8) $\mathcal{L}_{\text{kl}}=\mathbb{E}_{(X^{a},X^{t})\sim\mathcal{D}_{k}}\left[\text{KL}\left(E_{a}\parallel E_{t}\right)+\text{KL}\left(E_{t}\parallel E_{a}\right)\right].$
We propose two additional losses to address incremental learning tasks. To mitigate the forgetting issue in incremental learning and ensure consistency and alignment of the model across different feature spaces, we apply the losses $\mathcal{L}_{FD}$ and $\mathcal{L}_{SD}$ defined in Eq. 3 and Eq. 5.
@@
(9) $\mathcal{L}=\mathcal{L}_{\text{kl}}+\lambda\mathcal{L}_{\text{contrast}}+\mathcal{L}_{\text{FD}}+\alpha\mathcal{L}_{\text{SD}},$
@@
4. Experiments
4.1. Experimental Setup
Datasets. We conduct experiments with our PTAT framework, comparing it to baselines on the AudioCaps (Kim et al., 2019), Clotho (Drossos et al., 2020), AudioSet Strongly-labelled (hereafter referred to as AudioSet), and BBC Sound Effects subsets within the WavCaps (Mei et al., 2024) dataset. AudioCaps (Kim et al., 2019) contains approximately 50,000 audio clips sourced from AudioSet (Gemmeke et al., 2017), the largest audio event dataset, and is annotated by human listeners. Clotho (Drossos et al., 2020) comprises around 6,000 audio clips from the FreeSound2 platform, each clip accompanied by five human-annotated captions. The AudioSet and BBC Sound Effects subsets are part of the WavCaps (Mei et al., 2024) dataset.
Since neither AudioSet nor BBC Sound Effects provides predefined training and test splits, we manually divided the data for these datasets. The AudioSet subset contains 58,000 audio clips (after excluding portions overlapping with AudioCaps from the original AudioSet), which we randomly split into a training set of 55,000 samples and a test set of 3,000 samples. The BBC Sound Effects dataset consists of more than 30,000 audio clips recorded worldwide over the past century, which we randomly divided into a training set of 28,000 samples and a test set of 2,000 samples.
For our incremental learning setup, we define four incremental steps: AudioCaps as the first step, Clotho as the second, BBC Sound Effects as the third, and AudioSet as the fourth. These datasets vary in both content and structure. AudioCaps places greater emphasis on human speech and voice. In contrast, Clotho features longer audio clips and emphasizes diverse environmental sounds more. BBC Sound Effect focuses primarily on sound effects, such as those used in television and media production. AudioSet covers the broadest range of sound categories, with the highest diversity and complexity in audio content. These distinctions introduce domain gaps, making them well-suited for evaluating generalization and forgetting in incremental learning settings.
Table 1. Comparison with Baselines: Results on each dataset after sequential training across four datasets (AudioCaps, Clotho, BBC Sound Effects, and AudioSet). The results present the model’s performance at the Current Step of the experiment. Bold and underline indicate the best and the second-best results.
Method Parameters AudioCaps (Introduced at Step 1) Clotho (Introduced at Step 2)